It takes joe to fill the pool with a bucket in 30minutes. It take sue to fill the same pool in 45 minutes, and it takes bill to fill the same pool in 90minutes. How long would it take for all three to fill the pool at the same time?
Answer by ptaylor(2198)

Let x=time it takes all three working together to fill the pool.
Then all three work at the rate of 1/x pool per min
Joe works at the rate of 1/30 pool per min
Sue works at the rate of 1/45 pool per min
Bill works at the rate of 1/90 pool per min
Together they work at the rate of 1/30+1/45+1/90=
3/90 +2/90 +1/90=6/90=1/15 pool per min
Soooooo our equation to solve is:
1/15=1/x multiply each side by 15x
x=15 min ---time it takes all three working together
CK
In 15 min Joe completes (1/30)*15=1/2 of the pool
In 15 min Sue completes (1/45)*15=1/3 of the pool
In 15 min Bill completes (1/90)*15=1/6 of the pool
1/2 + 1/3 +1/6=1/2+3/6 =1 (pool that is)
Hope this helps---ptaylor